OpenStudy (anonymous):

Simplify the expression -5+i/2i Please explain

OpenStudy (anonymous):

does "simplify" mean "write in standard form"?

OpenStudy (anonymous):

you can multiply the top and bottom by \(i\) and get \[\frac{-5i-1}{-2}\] as a first step

OpenStudy (anonymous):

yes i think so...

OpenStudy (anonymous):

divide each term by \(-2\) get \[\frac{1}{2}+\frac{5}{2}i\] that is standard form

OpenStudy (anonymous):

thank you!!!